# Surface solar radiation: AI satellite retrieval can outperform Heliosat and generalizes well to other climate zones

### v1

@@ 3.3 Grid data preprocessing
The research domain of our study was selected as 29° N–62° N and 9° W–28° E to include most of Europe and parts of Northern Africa. Given the availability of high-quality measurements from IEA-PVPS in the years 2016-2022, we chose SARAH-3 and SEVIRI images from 2016 to 2023 for this study. The SEVIRI Level-1.5 data was acquired from the EUMETSAT datastore. We utilized eleven visible and infrared SEVIRI channels. The SEVIRI Level-1.5 data corresponds to image data that is corrected for unwanted radiometric and geometric effects. The Level-1.5 VIS006, VIS008, and IR016 reflectivity radiances were converted to bidirectional reflectance factor. The Level-1.5 emissivity radiances for channels IR039, WV062, WV073, IR087, IR097, IR108, IR120 and IR134 were converted to brightness temperature using the Satpy module [37]. The conversions homogenized the record between different MSG satellites and corrected the reflectivity channels for changing Sun-Earth distance. The images were reprojected to the geographic SARAH-3 grid with the nearest neighborhood interpolation.
@@ 3.4 Ground stations
Our SSI emulator models were validated on ground-based SSI sensors. Pyranometers measure the total amount of solar radiation from a half-dome in segments of one to ten minutes. These sparse ground observations represent only a single point in space whereas the SEVIRI imager measures an average reflectivity over a ground area of approximately 3x3 km. The ground station datasets for validating the SSI emulators are summarised in Table 1. They include SSI measurements from the Royal Netherlands Meteorological Institute (KNMI), the German Weather Service (DWD) and the Swiss Federal Office for Meteorology and Climatology (MeteoSwiss). Additionally, SSI measurements from the International Energy Agency Photovoltaic Power Systems Programme (IEA-PVPS) worldwide benchmark were gathered. They contain quality controlled SSI measurements from across Europe and North Africa [38]. The IEA-PVPS dataset comprises 14 stations from different operators. Six stations – CAB, CAR, CEN, PAL, PAY, and TOR – form part of the Baseline Surface Radiation Network (BSRN, [39]). Two stations – TAB and MIL – belong to the Centre for Energy, Environmental and Technological Research (CIEMAT) and Research on Energy Systems (RSE), respectively. The Swedish Meteorological and Hydrological Institute (SMHI) provided two additional stations, NOR and VIS. Finally, the EnerMENA project contributed four stations located in Northern Africa: GHA, MIS, OUJ, and TAT [40]. All stations of the IEA-PVPS dataset are equipped with three thermophile radiometers measuring SSI, direct normal irradiance and diffuse irradiance at 1-minute intervals [38].
@@
Dataset | Quality control | Stations | Measurement frequency (minutes) | Years |
IEA-PVPS | Yes | 14 | 1 | 2016-2022 |
DWD | Automatic | 99 | 10 | 2016-2022 |
KNMI | No | 34 | 10 | 2016-2022 |
MeteoSwiss | Automatic | 135 | 10 | 2016-2022 |
@@
The DWD dataset comprises SSI measurements from 99 radiometers including 70 scanning pyrheliometers (ScaPP) and 19 Kipp & Zonen CM11 or CM21 pyranometers [41]. The ScaPP instruments operate within a wavelength range of 0.3 to 1.1 based on a silicon detector. The MeteoSwiss dataset includes SSI measurements from 135 weather stations across Switzerland, all equipped with Kipp & Zonen CM21 pyranometers, with multiple stations located in Alpine regions. The KNMI dataset consists of 34 automatic weather stations, each featuring a single radiometer. This dataset is provided without quality control. All measurement stations are depicted in Figure 1.
@@
Due to its extensive spatial coverage and stringent quality control, the IEA-PVPS dataset served as the reference dataset for validating models and comparing their performance. The other datasets receive less weight in the analysis. The exact date ranges of observations may vary across ground stations to some extent for each of the four datasets (Table 1) because some stations were commissioned or decommissioned during the timeframe. Filtering on quality control can leave significant gaps in the data record as well.
@@ 4 SSI retrieval model
We trained five retrieval models in total. All models were trained on the same predictors: The Meteosat visible and infrared channels and the above scalar variables. The models are: E) a model trained only on DWD ground station data as target variable (section 4.2); A) an emulator model trained on the Heliosat SARAH-3 SSI as surrogate (section 4.3); and B-D) Three fine-tuned models, which trained to emulate SARAH-3 first and then fine-tuned on ground station data (section 4.4). The models are summarized in Table 2. An example case of SSI retrieved across the entire domain is shown in Figure 2 for Heliosat and all five models.
@@ 4.1 Architecture
We implemented a convolutional residual network (ConvResNet) similar to the one of Jiang et al. [23] with the following adjustments, as shown in Figure 9 in the Appendix. We defined patches of 15x15 pixels so that the centre pixel corresponds to the retrieved SSI pixel using all 11 visible and infrared channels. Extra geographical information was given by the digital elevation model (DEM) from Copernicus GLO-90 [42] coarsened to the same grid. Channel features were extracted from the patches by the CNN. Further features included the day of the year, the latitude, longitude, solar zenith angle and solar azimuth angle and were appended to the channel features and fed through a two-layer fully-connected neural network (FCN). The first part of the CNN is the feature mapping part where the spatial size is kept constant within the first two convolutional mappings (Figure 9). The second part consists of a 2x2 max pooling (MP) step to reduce spatial size followed by a convolutional residual block. The third part is the same where a 2x2 MP is followed by a convolutional residual block. The output of the CNN is first global average pooled (GAP) and then flattened to serve as input to the FCN. The kernel size in each convolutional layer is 3x3. Each convolutional layer in the network is followed by a 2D batch normalisation and ReLU activation function. Each of the two FCN layers is followed by a 1D batch normalisation and ReLU activation. The scalar output for the DL network is the SSI and was clipped for validation and inference to zero when estimating negative SSI. The input features SZA and solar azimuth angle (AZI) were computed at DEM altitude and nominal capture time. The retrieval with ConvResNet takes 15 seconds to run on a single Nvidia Tesla P100 GPU for a domain size of Europe (658x736 pixels). The retrieval could be linearly parallelized for larger domains or lower inference times. A newer and faster GPU than the Nvidia Tesla P100 (2016) would speed up the inference as well.
@@ 4.2 Training on ground stations
Colocated samples in time and space were created between the ground station and images to train a retrieval model on ground stations. The IEA-PVPS ground observations were averaged to 10 minutes from the start of each nominal satellite capture time. For the ground observations from KNMI, DWD and MeteoSwiss which already measured 10-minute averages, the SEVIRI nominal capture times were matched with ground observation measurement start times. This meant that non-quarterly observation times at the 10/20/40/50 minute marks were disregarded. All available matching observations between 2016 and 2022 were used. Patches of 15x15 pixels with the center pixel overlapping the ground station were selected. Scalar input features latitude, longitude, SZA and AZI were calculated for the exact locations of the respective stations. The DWD stations were split randomly into 80 training and 19 validation stations. The density of the MeteoSwiss network is so high that we could afford to keep more validation stations than training stations and they were randomly split into 50 training and 85 validation stations. The KNMI stations were left out of training due to the lack of quality control. The station data were filtered for SZA¡ to select daytime periods only. When training on ground stations, the samples of different sites were first merged and then randomly shuffled to ensure that training, validation and test sets have similar statistical characteristics. The model (E) was trained only on the DWD dataset for 15 epochs with a mean squared error loss with an Adam optimizer with weight decay [43].
@@ 4.3 Emulator training on Heliosat
The emulator model (A) was trained on Heliosat SARAH-3 SSI. SARAH-3 provides SSI fields for every second capture of the SEVIRI imager. The SARAH-3 SSI as target variable and the SEVIRI images and scalar features (latitude, longitude, SZA, AZI, day of the year) as predictor variables were merged at nominal capture time to form a dataset to train, validate and test our SSI emulator model. The scalar features were calculated at the center of each pixel. We split the SEVIRI images and SARAH-3 retrievals into training (years 2016–2021), validation (2022) and test set (2023). Daytime locations were defined as pixels where SZA . Images with less than 50% daytime pixels were dropped. The images were shuffled randomly for training. A sample consists of 15x15 patches of each grid input and the five scalar features from the center pixel. The inputs and outputs were min-max normalized for all training and fine-tuning steps. A training batch consisted of 2048 uniformly distributed samples per image. Only the daytime pixels were sampled from images containing both day- and nighttime pixels. A single epoch consisted of cycling through all images with 2048 samples per image. The network was trained for two epochs on a mean squared error loss with an Adam optimizer with weight decay [43]. The validation loss was monitored to prevent overfitting the training dataset. A learning rate of dropping to in the second epoch yielded a balance between training time and recognizing of overfitting. Every 10% of a full training epoch, the equivalent of half a year of SEVIRI captures, the validation metrics were calculated and a checkpoint was made. From those checkpoints, the best model was picked.
@@ 4.4 Fine-tuning on ground stations
In total, we trained three fine-tuned models (rows B, C and D in Table 2). To arrive at the first fine-tuned retrieval model (B), we fine-tuned the emulator on the DWD ground stations training set with a proximity loss (), a frozen FCN and a decreased learning rate . In the second stage of fine-tuning, which yielded the second fine-tuned model (C), the previous model fine-tuned on DWD (B) was further fine-tuned on the MeteoSwiss training set with a higher proximity loss (), a frozen FCN and the same learning rate. In the last stage of fine-tuning, which yielded the third fine-tuned model (D), the model previously fine-tuned on DWD and MeteoSwiss ground stations (C) was further fine-tuned on the IEA-PVPS SSI measurements without a frozen FCN and with a smaller learning rate lr. For each stage, the previous model weights were set as the in the proximity loss, as shown in column ”Weights ” in Table 2. In the first two stages of fine-tuning, the validation loss and MBE per station on IEA-PVPS were monitored ten times per epoch to signalize a change in bias or overfitting to the stations on which the model was being fine-tuned. In the third stage, the validation loss and MBE on MeteoSwiss and DWD validation stations were monitored for the same reason. Each fine-tuning stage lasted at most two epochs on their respective training set or shorter when the model showed signs of overfitting or bias shifts.
Models | Target data seen | Learning rate |
FCN freeze | ||||||||||||
@@ 5.1 Characterisation of the SSI emulator
The emulator model trained in section 4.2 was compared to the Heliosat SARAH-3 SSI on the test set and to the ground stations SSI to characterise its performance. Root mean square deviation (RMSD) and mean bias deviation (MBD) were calculated pixel-wise and over daytime periods only. We found that the emulator follows the SARAH-3 SSI remarkably closely, as shown in Figures 3 and 4. It emulates SSI with high accuracy and with biases negligibly small for solar energy applications such as resource assessments and forecasting. Specifically, the emulator model has a small positive MBD of 10.2 by somewhat overestimating low Heliosat SARAH-3 SSI values (Figure 3). The emulator generates an RMSD of 51.7 with regard to the SARAH-3 SSI on the test set.
The RMSD with regard to SARAH-3 is the largest in mountainous regions such as the Alps, the Pyrenees, and the Scandinavian mountains, with RMSD values of up to 125 . Stations in mountainous regions tend to be affected by larger SSI retrieval uncertainties due to complex terrain, local clouds and snow cover [2]. Snow cover can be challenging to distinguish from clouds due to the similarity of visible reflectivity, which leads to underestimation when snow cover is mistaken for clouds [2]. Non-moving highly reflective pixels are flagged and treated either as fog or snow cover by Heliosat SARAH-3 [28]. This inadvertently can lead to the opposite problem of overestimating SSI where it accurately identifies snow cover but does not consider clouds covering a snowy region.
Deviations with regard to Heliosat SSI are also somewhat elevated across North Africa with RMSD values of around 75 . The lowest regional RMSD of around 45 can be found over large water bodies, such as the North Sea, the Atlantic and the Mediterranean Sea. The common denominator for elevated RMSD is the persistent surface albedo during the year. Snow cover but also the bare surfaces of North African Sahara show increased reflectivity most of the year [44]. Water bodies on the contrary show a low amount of backscatter for most incidence angles and thus have generally low albedo [45]. A high uncertainty of the Heliosat method comes from large clear-sky reflectances [28]. Large clear-sky reflectances lower the contrast between clear-sky and cloudy reflectance which linearly impacts the uncertainty in CSI and subsequently SSI of SARAH-3. The heightened RMSD in these regions might be mostly due to the ground truth being noisy or inaccurate.

### v2

@@ 3.3 Grid data preprocessing
The research domain of our study was selected as 29° N–62° N and 9° W–28° E to include most of Europe and parts of Northern Africa. Given the availability of high-quality measurements from IEA-PVPS in the years 2016-2022, we chose SARAH-3 and SEVIRI images from 2016 to 2023 for this study. The SEVIRI Level-1.5 data was acquired from the EUMETSAT datastore. We utilized eleven visible and infrared SEVIRI channels. The SEVIRI Level-1.5 data corresponds to image data that is corrected for unwanted radiometric and geometric effects. The Level-1.5 VIS006, VIS008, and IR016 reflectivity radiances were converted to bidirectional reflectance factor. The Level-1.5 emissivity radiances for channels IR039, WV062, WV073, IR087, IR097, IR108, IR120 and IR134 were converted to brightness temperature using the Satpy module (Raspaud et al. 2024). The conversions homogenized the record between different MSG satellites and corrected the reflectivity channels for changing Sun-Earth distance. The images were reprojected to the geographic SARAH-3 grid with the nearest neighborhood interpolation.
@@ 3.4 Ground stations
Our SSI emulator models were validated on ground-based SSI sensors. Pyranometers measure the total amount of solar radiation from a half-dome in segments of one to ten minutes. These sparse ground observations represent only a single point in space whereas the SEVIRI imager measures an average reflectivity over a ground area of approximately 33 km. The ground station datasets for validating the SSI emulators are summarised in Table 3.4. They include SSI measurements from the Royal Netherlands Meteorological Institute (KNMI), the German Weather Service (DWD) and the Swiss Federal Office for Meteorology and Climatology (MeteoSwiss). Additionally, SSI measurements from the International Energy Agency Photovoltaic Power Systems Programme (IEA-PVPS) worldwide benchmark were gathered. They contain quality controlled SSI measurements from across Europe and North Africa (Forstinger et al. 2022). The IEA-PVPS dataset comprises 14 stations from different operators. Six stations – CAB, CAR, CEN, PAL, PAY, and TOR – form part of the Baseline Surface Radiation Network (BSRN, (Driemel et al. 2018)). Two stations – TAB and MIL – belong to the Centre for Energy, Environmental and Technological Research (CIEMAT) and Research on Energy Systems (RSE), respectively. The Swedish Meteorological and Hydrological Institute (SMHI) provided two additional stations, NOR and VIS. Finally, the EnerMENA project contributed four stations located in Northern Africa: GHA, MIS, OUJ, and TAT (Schüler et al. 2016). All stations of the IEA-PVPS dataset are equipped with three thermophile radiometers measuring SSI, direct normal irradiance and diffuse irradiance at 1-minute intervals (Forstinger et al. 2022).
@@
Datasets of SSI ground station measurements. All years are included in the training/validation. Details Dataset Quality control Stations Measurement frequency (minutes) Years IEA-PVPS Yes 14 1 2016-2022 DWD Automatic 99 10 2016-2022 KNMI No 34 10 2016-2022 MeteoSwiss Automatic 135 10 2016-2022
@@
The DWD dataset comprises SSI measurements from 99 radiometers including 70 scanning pyrheliometers (ScaPP) and 19 Kipp & Zonen CM11 or CM21 pyranometers (Becker and Behrens 2012). The ScaPP instruments operate within a wavelength range of 0.3 to 1.1 based on a silicon detector. The MeteoSwiss dataset includes SSI measurements from 135 weather stations across Switzerland, all equipped with Kipp & Zonen CM21 pyranometers, with multiple stations located in Alpine regions. The KNMI dataset consists of 34 automatic weather stations, each featuring a single radiometer. This dataset is provided without quality control. All measurement stations are depicted in Figure 1.
@@
Due to its extensive spatial coverage and stringent quality control, the IEA-PVPS dataset served as the reference dataset for validating models and comparing their performance. The other datasets receive less weight in the analysis. The exact date ranges of observations may vary across ground stations to some extent for each of the four datasets (Table 3.4) because some stations were commissioned or decommissioned during the timeframe. Filtering on quality control can leave significant gaps in the data record as well.
@@ 4 SSI retrieval model
We trained five retrieval models in total. All models were trained on the same predictors: The Meteosat visible and infrared channels and the above scalar variables. The models are: E) a model trained only on DWD ground station data as target variable (section 4.2); A) an emulator model trained on the Heliosat SARAH-3 SSI as surrogate (section 4.3); and B-D) Three fine-tuned models, which trained to emulate SARAH-3 first and then fine-tuned on ground station data (section 4.4). The models are summarized in Table 4.4. An example case of SSI retrieved across the entire domain is shown in Figure 2 for Heliosat and all five models.
@@ 4.1 Architecture
We implemented a convolutional residual network (ConvResNet) similar to the one of Jiang et al. (2019) with the following adjustments, as shown in Figure 9 in the Appendix. Instead of 1616 we defined patches of 1515 pixels so that the centre pixel corresponds to the retrieved SSI pixel using all 11 visible and infrared channels. The output of our network is instantaneous SSI instead of hourly averaged SSI. Jiang et al. (2019) argued against including the solar zenith and azimuth angles in the network, suggesting that the spatial information is inherently captured by the convolutions. However, we find no compelling reason to exclude these variables, as they are strongly correlated with SSI and can be easily precomputed. What’s more, is that we applied batch normalization between each convolutional layer to prevent overfitting of the model. Lastly, extra geographical information was given by the digital elevation model (DEM) from Copernicus GLO-90 (AIRBUS 2024) coarsened to the same grid as MSG.
The first part of the CNN is the feature mapping part where the spatial size is kept constant within the first two convolutional mappings (Figure 9). Channel features were extracted from the patches by the CNN. Further features included the day of the year, the latitude, longitude, solar zenith angle and solar azimuth angle and were appended to the channel features and fed through a two-layer fully-connected neural network (FCN). The second part consists of a 22 max pooling (MP) step to reduce spatial size followed by a convolutional residual block. The third part is the same where a 22 MP is followed by a convolutional residual block. The output of the CNN is first global average pooled (GAP) and then flattened to serve as input to the FCN. The kernel size in each convolutional layer is 33. Each convolutional layer in the network is followed by a 2D batch normalisation and ReLU activation function. Each of the two FCN layers is followed by a 1D batch normalisation and ReLU activation.
The scalar output for the DL network is the SSI and was clipped for validation and inference to zero when estimating negative SSI. The input features SZA and solar azimuth angle (AZI) were computed at DEM altitude and nominal capture time. The retrieval with ConvResNet takes 15 seconds to run on a single Nvidia Tesla P100 GPU for a domain size of Europe (658736 pixels). The retrieval could be linearly parallelized for larger domains or lower inference times. A newer and faster GPU than the Nvidia Tesla P100 (2016) would speed up the inference as well.
@@ 4.2 Training on ground stations
Colocated samples in time and space were created between the ground station and images to train a retrieval model on ground stations. The IEA-PVPS ground observations were averaged to 10 minutes from the start of each nominal satellite capture time. For the ground observations from KNMI, DWD and MeteoSwiss which already measured 10-minute averages, the SEVIRI nominal capture times were matched with ground observation measurement start times. This meant that non-quarterly observation times at the 10/20/40/50 minute marks were disregarded. All available matching observations between 2016 and 2022 were used. Patches of 1515 pixels with the center pixel overlapping the ground station were selected. Scalar input features latitude, longitude, SZA and AZI were calculated for the exact locations of the respective stations. The DWD stations were split randomly into 80 training and 19 validation stations. The density of the MeteoSwiss network is so high that we could afford to keep more validation stations than training stations, and they were randomly split into 50 training and 85 validation stations. The KNMI stations were left out of training due to the lack of quality control. The station data were filtered for SZA to select daytime periods only. When training on ground stations, the samples of different sites were first merged and then randomly shuffled to ensure that training, validation and test sets have similar statistical characteristics. The model (E) was trained only on the DWD dataset for 15 epochs with a mean squared error loss with an Adam optimizer with weight decay (Kingma and Ba 2014).
@@ 4.3 Emulator training on Heliosat
The emulator model (A) was trained on Heliosat SARAH-3 SSI. SARAH-3 provides SSI fields for every second capture of the SEVIRI imager. The SARAH-3 SSI as target variable and the SEVIRI images and scalar features (latitude, longitude, SZA, AZI, day of the year) as predictor variables were merged at nominal capture time to form a dataset to train, validate and test our SSI emulator model. The scalar features were calculated at the center of each pixel. We split the SEVIRI images and SARAH-3 retrievals into training (years 2016–2021), validation (2022) and test set (2023). Daytime locations were defined as pixels where SZA . Images with less than 50% daytime pixels were dropped. The images were shuffled randomly for training. A sample consists of 1515 patches of each grid input and the five scalar features from the center pixel. The inputs and outputs were min-max normalized for all training and fine-tuning steps. A training batch consisted of 2048 uniformly distributed samples per image. Only the daytime pixels were sampled from images containing both day- and nighttime pixels. A single epoch consisted of cycling through all images with 2048 samples per image. The network was trained for two epochs on a mean squared error loss with an Adam optimizer with weight decay (Kingma and Ba 2014). The validation loss was monitored to prevent overfitting the training dataset. A learning rate of dropping to in the second epoch yielded a balance between training time and recognizing of overfitting. Every 10% of a full training epoch, the equivalent of half a year of SEVIRI captures, the validation metrics were calculated and a checkpoint was made. From those checkpoints, the best model was picked.
@@ 4.4 Fine-tuning on ground stations
In total, we trained three fine-tuned models (rows B, C and D in Table 4.4). To arrive at the first fine-tuned retrieval model (B), we fine-tuned the emulator on the DWD ground stations training set with a proximity loss (), a frozen FCN and a decreased learning rate . In the second stage of fine-tuning, which yielded the second fine-tuned model (C), the previous model fine-tuned on DWD (B) was further fine-tuned on the MeteoSwiss training set with a higher proximity loss (), a frozen FCN and the same learning rate. In the last stage of fine-tuning, which yielded the third fine-tuned model (D), the model previously fine-tuned on DWD and MeteoSwiss ground stations (C) was further fine-tuned on the IEA-PVPS SSI measurements without a frozen FCN and with a smaller learning rate lr. For each stage, the previous model weights were set as the in the proximity loss, as shown in column ’Weights ’ in Table 4.4. In the first two stages of fine-tuning, the validation loss and MBE per station on IEA-PVPS were monitored ten times per epoch to signalize a change in bias or overfitting to the stations on which the model was being fine-tuned. In the third stage, the validation loss and MBE on MeteoSwiss and DWD validation stations were monitored for the same reason. Each fine-tuning stage lasted at most two epochs on their respective training set or shorter when the model showed signs of overfitting or bias shifts.
Specification of the different trained models.
| Target data seen | Learning rate |
| Models |
@@ 5.1 Characterisation of the SSI emulator
The emulator model trained in section 4.2 was compared to the Heliosat SARAH-3 SSI on the test set and to the ground stations SSI to characterise its performance. Instantaneous root-mean-square deviation (RMSD) and mean bias deviation (MBD) were only calculated pixel-wise and over daytime periods. We found that the emulator follows the SARAH-3 SSI remarkably closely, as shown in Figures 3 and 4. It emulates SSI with high accuracy and with biases negligibly small for solar energy applications such as resource assessments and forecasting. Specifically, the emulator model has a small positive MBD of 10.2 by somewhat overestimating low Heliosat SARAH-3 SSI values (Figure 3). The emulator generates an RMSD of 51.7 with regard to the SARAH-3 SSI on the test set.
The RMSD with regard to SARAH-3 is the largest in mountainous regions such as the Alps, the Pyrenees, and the Scandinavian mountains, with RMSD values of up to 125 . Stations in mountainous regions tend to be affected by larger SSI retrieval uncertainties due to complex terrain, local clouds and snow cover (Carpentieri et al. 2023). Snow cover can be challenging to distinguish from clouds due to the similarity of visible reflectivity, which leads to underestimation when snow cover is mistaken for clouds (Carpentieri et al. 2023). Non-moving highly reflective pixels are flagged and treated either as fog or snow cover by Heliosat SARAH-3 (Pfeifroth et al. 2023). This inadvertently can lead to the opposite problem of overestimating SSI where it accurately identifies snow cover but does not consider clouds covering a snowy region.
Deviations with regard to Heliosat SSI are also somewhat elevated across North Africa with RMSD values of around 75 . The lowest regional RMSD of around 45 can be found over large water bodies, such as the North Sea, the Atlantic and the Mediterranean Sea. The common denominator for elevated RMSD is the persistent surface albedo during the year. Snow cover but also the bare surfaces of North African Sahara show increased reflectivity most of the year (Bierwirth et al. 2009). Water bodies on the contrary show a low amount of backscatter for most incidence angles and thus have generally low albedo (Jin et al. 2011). A high uncertainty of the Heliosat method comes from large clear-sky reflectances (Pfeifroth et al. 2023). Large clear-sky reflectances lower the contrast between clear-sky and cloudy reflectance which linearly impacts the uncertainty in CSI and subsequently SSI of SARAH-3. The heightened RMSD in these regions might be mostly due to the ground truth being noisy or inaccurate.